# references.bib

### v2

@@ 4.1 Task
The task used in this research is the spatial navigation problem of navigating an unknown maze to reach a goal whose location is not known in advance. At each location, the robot has eight possible actions, each varying in its directions. A learning episode starts with the robot at the starting point set depending on the maze. At every step, the agent will choose an action and will receive a corresponding reward for its choice. This is repeated until the goal is reached or the number of steps has reached the set horizon to prevent an infinite amount of time in completing the learning task. The reward structure is critical in reinforcement learning, and hierarchical reinforcement learning is no exception. For this problem the reward structure is a large penalty for hitting a wall, a reward for reaching the goal, and a small penalty for any other state.
@@ 4.2 Simulation Environment
For the environment in these experiments, the Webots simulator was chosen along with the open-source FAIRIS framework to assist in robot control [fairis]. An example of the Webots simulator is shown in Figure 6. This was chosen to simulate the real world as closely as possible when testing the algorithm. The University of South Florida’s GAIVI supercomputer was used to perform the experiments, which on average took three days to complete.
@@ 4.4 Exploration Structure
For the Proximal Policy Optimization [PPO], the exploration structure was set similar to the Option Critic Architecture. This structure was to randomly sample from the distribution created by the actor. As this was found not to explore enough and not to converge due to insufficient exploration, a version of the decaying $\epsilon$-greedy algorithm was used instead. This version was to sample a uniform distribution and if less than $\epsilon$ then a random action is chosen and $\epsilon$ is decayed. Otherwise, an action is chosen according to the PPO policy.
@@ 5.1 Mazes
One of the standard benchmarks for hierarchical reinforcement learning is the four-rooms environment from Sutton et al. [SUTTON]. In the experiments to compare OC with PPO, two open mazes with obstacles were used (one-room multiple small obstacles maze and one-room single large obstacle maze) and the four-rooms experiments, as shown in Figure 7. The maze in Figure 7(a) is simple as there is only one room, and the obstacles are relatively small and randomly placed. The second maze in Figure 7(b) is harder to solve, as it has one large obstacle that is placed directly between the agent and the goal, where the agent has to learn to go further away from the goal to be able to go around the obstacle and reach the goal. The third maze in Figure 7(c) is the most difficult, corresponding to the four-rooms maze.
@@ 5.3 Experiment 2: Difference Between Critic and Termination Automatically Created Sub-Goals
There are many different ways to automatically create the sub-goals. In this experiment, we compare critic created sub-goals and termination created sub-goals. Critic created sub-goals are where the critic repeatedly chooses the same option until reaching a certain sub-goal state. Although the policy over options chooses which option to use, the critic directly influences this decision, which is why they are called critic sub-goals. Termination sub-goals are the state where the termination function decides to end the current option. It is hypothesized that the termination created sub-goals will perform better than the critic created sub-goals because as shown in Equation 5 the termination function is updated to compare the differences between options to evaluate when to switch when in contrast as shown in Equation 2 the critic is updated to evaluate an option in the current state. The original OC model uses termination created sub-goals [OptionCritic]. To compare to the critic created sub-goals, the termination function has been removed completely, and instead, the option will be terminated at every step. This is done to remove the termination function and prevent it from creating sub-goals, so the critic will set the sub-goals instead. This experiment will be tested on the four-rooms maze. This modified algorithm was then tested and compared to the original algorithm the results are shown in Table 6 and Figure 9. A student’s T-test was used on both the convergence time and path lengths to see if the differences were statistically significant, also shown in Table 6.
@@ 5.4 Experiment 3: Comparison Between Automatic and Manual Sub-Goals
This experiment is designed to test the differences between automatically set sub-goals and manual set sub-goals. In this experiment, the sub-goals will be created by the termination function and will be tested on the four-rooms maze. It is hypothesized that the algorithm will perform better if it is able to create the sub-goals instead of them being set by hand. This is because the algorithm can set sub-goals to help explore at the beginning of the training. To test this hypothesis, the algorithm was modified to have the sub-goals set by hand. To do this, the termination code was removed and replaced with code checking if the algorithm was in a specific state. Then, the states to terminate were set by hand with sub-goals that divide the task into equal parts to make each sub-task equal in length. Each algorithm was run five times, and a student’s T-test was performed on both the path length and convergence time shown in Table 7. This table shows that the model set sub-goals convergence about twice as fast, and the paths are about ten steps shorter.
@@ 5.5 Experiment 4: Performance Evaluation on Frequency of Terminations
One of the expected strengths of applying Hierarchical Reinforcement Learning is its ability to break down tasks into sub-tasks and create sub-goals. In this experiment termination created sub-goals will be used. Sub-goals are an important part of the hierarchy created by the algorithm because they define how the task is divided into parts. Because of it’s importance this experiment tests the effects of the frequency of termination on the performance of the algorithm. The frequency of termination determines how many sub-goals are created, as the sub-goal is the state an option reaches before switching to the next option. Additionally if the termination function does not create enough sub-goals then the problem won’t be split up correctly and one option could end up covering the whole task, also known as degerenating. The termination regulation, $\phi$ value was added to the termination function update equation to control the frequency of termination as shown in Equation 2. It is hypothesized as the $\phi$ shrinks, the algorithm will be able to create better termination sub-goals because the smaller $\phi$ will prevent the options from growing too large and degenerating. To test this hypothesis, the Option Critic will be run on the four-rooms maze while varying the $\phi$. All the hyperparameters will be fixed while the $\phi$ is varied. $\phi$ will be varied between 0 and 0.1 with increments of 0.01. These values were chosen to show how the performance changes as the options grow in size. In the original model, the value of 0.01 was used to incentivize creating larger options. Values larger than 0.01 were chosen to test whether incentivizing the creation of larger options is necessary. This experiment will be performed on the four-rooms maze. The maximum number of steps used in this experiment is 500000 steps. This experiment was run five times to verify the results, and so a statistical test could be performed. An ANOVA test was used to test the differences between each termination value for each of the metrics: convergence time, option lengths, and path length. The results of this experiment are shown in Table 8. To better understand the results, several graphs of one of the experiments are constructed and shown below in Figures 10, 11, and 12. In Figure 10, the option lengths are shown for each of the termination regulation values. In Figure 11, the path lengths are shown as the algorithm is trained for each of the tested values. In Figure 12, the shortest path is shown for each of the test values.
@@ 6 Conclusions and Discussion
Hierarchical Reinforcement Learning is a powerful tool that can solve complex problems that standard reinforcement learning struggles with. HRL is able to build a hierarchy that breaks down tasks into smaller sub-tasks by defining sub-goals. These sub-goals may have a large impact on the performance of the algorithm and solve complicated problems. This research has four contributions to the scientific community: a comparison of standard reinforcement learning to hierarchical reinforcement learning, an evaluation of two different ways to automatically create sub-goals, a comparison of manual and automatically created sub-goals, and a performance evaluation of the frequency of termination.
@@ 6.1 Experiment 1: Comparison Between OC and PPO
OC was shown to perform better when compared to PPO as the complexity increases. The results show that with one-room multiple small obstacles maze, the PPO has a similar convergence performance to the OC, and as the maze complexity increases, the convergence performance of PPO compared to OC decreases steadily. This is shown in the Table 3 as the P-value for the one-room multiple small obstacles maze is 0.45, suggesting that the convergence performance is about the same. When the P-value for the one-room single large obstacle maze is less than 0.05 at 0.048, it is statistically significant, showing that the OC performs better than the PPO in terms of convergence time. With the four-rooms maze the difference is even larger where the OC converges about twice as fast as the PPO and has a P-value of 0.036. Although the convergence time is different between the PPO and OC, the path length is about the same. This was expected as they were both given time to converge, and they didn’t have difficulties finding a near optimal maze. Based on these results, a T-test was not performed on the path lengths.
@@ 6.3 Experiment 3: Comparison Between Automatic and Manual Sub-Goals
The idea behind sub-goals is to break down tasks into smaller tasks, such that the smaller tasks are easier for the policies to solve as they are not as complicated. But it is important that the task is split evenly. If not, one of the tasks will be more difficult and would remove the purpose of creating the smaller tasks to make easier tasks. Since it is important to break down the tasks correctly, it may seem best to set the sub-goals by hand to ensure the task is broken down correctly. To manually set the best possible sub-goals for this experiment, each doorway and the goal were chosen as the sub-goals. These were chosen because they evenly break down the task. This is similar to the originally designed four-rooms maze task, where each room was assigned a different option, with the doorway being the sub-goals [SUTTON]. As seen in Table 7 when the algorithm uses the manually set sub-goals it takes about twice as long to converge and it’s path length is about ten steps longer. These results support the hypothesis stating the algorithm will perform better if the sub-goals are set by the algorithm. These results support the hypothesis suggesting the algorithm is able to break down the tasks adequately, thus hand-set sub-goals are not required and even hinder the process as they impede exploration as most of the exploration in the Option Critic comes from trying different options and correspondingly pick a different action. The consideration is that if the sub-goals are set by hand, the algorithm wouldn’t be considering different options in the same state, explaining why automatic sub-goals produce better results. In summary the automatic sub-goals perform better because they allow the algorithm to explore more not because the sub-goals are necessarily better. In this research there was no metric to compare the sub-goals to ideal sub-goals but for future research such metric could be used to better understand the differences between sub-goals created by algorithm and sub-goals set by hand.
@@ 6.4 Experiment 4: Performance Evaluation on Frequency of Terminations
Theoretically, sub-goals are advantageous because they enable an algorithm to break down tasks into several smaller, simpler sub-tasks, which are easier for the algorithm to solve. These results support this hypothesis because the Option Critic is able to outperform the PPO as shown in Table 3. But Option Critics’ increased performance could come from other factors, in addition to sub-goals. In Table 8, it shows how as the termination regulation value increases from 0.00 to 0.10, the convergence time also increases. In Figure 11, it shows how as the termination regulation value increases, the sub-goals are longer and there are fewer sub-goals, up to the point where there are almost no sub-goals. In general, as the number of sub-goals decreases, the performance decreases, showing how it is important to have sub-goals. This might suggest that it would be best to terminate in every state, but as seen in experiment 3, this can also lead to poor performance. This shows how important sub-goals are in most situations, but this may not be true in all cases. As there may be simple problems that do not need to be broken up and thus do not require sub-goals. Furthermore, in simple problems, sub-goals may hinder performance by adding unnecessary complexity to the task.
@@ 6.5 Future Work
Future research will investigate further ways to set the sub-goals automatically. Currently, the policy over options relies mostly on the critic to create sub-goals, but this could be replaced by using a different algorithm for the policy over options, such as an additional network or a different machine learning technique. This could improve the exploration of the sub-goals and the overall algorithm performance. Future research will also investigate the performance of the Option Critic in physical robot learning tasks in spatial navigation tasks in the context of sequence learning [cazin] [medeiros].
[…]
[heading=bibintoc,title=References]

### v3

@@ 4.1 Task
The task used in this research is the spatial navigation problem of navigating an unknown maze to reach a goal whose location is not known in advance. At each location, the robot has eight possible actions, each varying in its direction. A learning episode starts with the robot at the starting point set depending on the maze. At every step, the agent will choose an action and will receive a corresponding reward for its choice. This is repeated until the goal is reached or the number of steps has reached the set horizon to prevent an infinite amount of time in completing the learning task. The reward structure is critical in reinforcement learning, and hierarchical reinforcement learning is no exception. For this problem the reward structure is a large penalty for hitting a wall, a reward for reaching the goal, and a small penalty for any other state.
@@ 4.2 Simulation Environment
For the environment in these experiments, the Webots simulator was chosen along with the open-source FAIRIS framework to assist in robot control fairis (6). An example of the Webots simulator is shown in Figure 6. This was chosen to simulate the real world as closely as possible when testing the algorithm. The University of South Florida’s GAIVI supercomputer was used to perform the experiments, which on average took three days to complete. Additionally, all of the code and data for this paper are located in FAIRIS fairis (6).
@@ 4.4 Exploration Structure
For the Proximal Policy Optimization PPO (34), the exploration structure was set similar to the Option Critic Architecture. This structure was to randomly sample from the distribution created by the actor. As this was found not to explore enough and not to converge due to insufficient exploration, a version of the decaying $\epsilon$-greedy algorithm was used instead. This version was to sample a uniform distribution and if less than $\epsilon$ then a random action is chosen and $\epsilon$ is decayed. Otherwise, an action is chosen according to the PPO policy.
@@ 5.1 Mazes
One of the standard benchmarks for hierarchical reinforcement learning is the four-rooms environment from Sutton et al. SUTTON (40). In the experiments to compare OC with PPO, two open mazes with obstacles were used (one-room multiple small obstacles maze and one-room single large obstacle maze) and the four-rooms experiments, as shown in Figure 7. The maze in Figure 7(a) is simple as there is only one room, and the obstacles are relatively small and randomly placed. The second maze in Figure 7(b) is harder to solve, as it has one large obstacle that is placed directly between the agent and the goal, where the agent has to learn to go further away from the goal to be able to go around the obstacle and reach the goal. The third maze in Figure 7(c) is the most difficult, corresponding to the four-rooms maze.
@@ 5.3 Experiment 2: Difference Between Critic and Termination Automatically Created Sub-Goals
There are many different ways to automatically create the sub-goals. In this experiment, we compare critic created sub-goals and termination created sub-goals. Critic created sub-goals are where the critic repeatedly chooses the same option until reaching a certain sub-goal state. Although the policy over options chooses which option to use, the critic directly influences this decision, which is why they are called critic sub-goals. Termination sub-goals are the state where the termination function decides to end the current option. It is hypothesized that the termination created sub-goals will perform better than the critic created sub-goals because as shown in Equation 5 the termination function is updated to compare the differences between options to evaluate when to switch when in contrast as shown in Equation 2 the critic is updated to evaluate an option in the current state. The original OC model uses termination created sub-goals OptionCritic (3). To compare to the critic created sub-goals, the termination function has been removed completely, and instead, the option will be terminated at every step. This is done to remove the termination function and prevent it from creating sub-goals, so the critic will set the sub-goals instead. This experiment will be tested on the four-rooms maze. This modified algorithm was then tested and compared to the original algorithm the results are shown in Table 6 and Figure 9. A student’s T-test was used on both the convergence time and path lengths to see if the differences were statistically significant, also shown in Table 6.
@@ 5.4 Experiment 3: Comparison Between Automatic and Manual Sub-Goals
This experiment is designed to test the differences between automatically set sub-goals and manually set sub-goals. In this experiment, the sub-goals will be created by the termination function and will be tested on the four-rooms maze. It is hypothesized that the algorithm will perform better if it is able to create the sub-goals instead of them being set by hand. This is because the algorithm can set sub-goals to help explore at the beginning of the training.
Figure 10: This figure shows the sub-goals for each option. The start point is shown in red, the goal in green, and the sub-goals in blue. Each option starts when the previous one ends.
To test this hypothesis, the algorithm was modified to have the sub-goals set by hand. To do this, the termination code was replaced with code that checks if the algorithm is in a specific handset state using a cosine similarity and terminates the current option accordingly. Then, the states to terminate were with sub-goals that divide the task into equal parts to make each sub-task equal in length. For each option, the sub-goal was set to the doorway to the next room or the goal if the goal was in the room. Figure 10 shows the sub-goals for each option. The doorways are used to break the problem down into four equal parts, dividing the work among the options. Although there are alternative ways to set the sub-goals, this was used because the four-rooms problem was designed to be easily broken down in this manner for testing hierarchical reinforcement learning algorithms SUTTON (40). Additionally, the option selection code was replaced to prevent the critic from choosing the wrong option for that room; instead, the option was selected according to the hand-set sub-goals. The doorways were chosen as the sub-goals because they even break up the problem as the four-room maze was designed to be broken down into rooms SUTTON (40).
Each algorithm was run five times, and a student’s T-test was performed on both the path length and convergence time shown in Table 7. This table shows that the model set sub-goals converge about twice as fast, and the paths are about ten steps shorter.
@@ 5.5 Experiment 4: Performance Evaluation on Frequency of Terminations
One of the expected strengths of applying Hierarchical Reinforcement Learning is its ability to break down tasks into sub-tasks and create sub-goals. In this experiment termination created sub-goals will be used. Sub-goals are an important part of the hierarchy created by the algorithm because they define how the task is divided into parts. Because of it’s importance this experiment tests the effects of the frequency of termination on the performance of the algorithm. The frequency of termination determines how many sub-goals are created, as the sub-goal is the state an option reaches before switching to the next option. Additionally if the termination function does not create enough sub-goals then the problem won’t be split up correctly and one option could end up covering the whole task, also known as degerenating. The termination regulation, $\phi$ value was added to the termination function update equation to control the frequency of termination as shown in Equation 2. It is hypothesized as the $\phi$ shrinks, the algorithm will be able to create better termination sub-goals because the smaller $\phi$ will prevent the options from growing too large and degenerating. To test this hypothesis, the Option Critic will be run on the four-rooms maze while varying the $\phi$. All the hyperparameters will be fixed while the $\phi$ is varied. $\phi$ will be varied between 0 and 0.1 with increments of 0.01. These values were chosen to show how the performance changes as the options grow in size. In the original model, the value of 0.01 was used to incentivize creating larger options. Values larger than 0.01 were chosen to test whether incentivizing the creation of larger options is necessary. This experiment will be performed on the four-rooms maze. The maximum number of steps used in this experiment is 500000 steps. This experiment was run five times to verify the results, and so a statistical test could be performed. An ANOVA test was used to test the differences between each termination value for each of the metrics: convergence time, option lengths, and path length. The results of this experiment are shown in Table 8. To better understand the results, several graphs of one of the experiments are constructed and shown below in Figures 11, 12, and 13. In Figure 11, the option lengths are shown for each of the termination regulation values. In Figure 12, the path lengths are shown as the algorithm is trained for each of the tested values. In Figure 13, the shortest path is shown for each of the test values.
@@ 6 Conclusions and Discussion
Hierarchical Reinforcement Learning is a powerful tool that can solve complex problems that standard reinforcement learning struggles with. HRL is able to build a hierarchy that breaks down tasks into smaller sub-tasks by defining sub-goals. These sub-goals may have a significant impact on the performance of the algorithm and solve complicated problems. The primary contribution of this research is to demonstrate the importance of sub-goals. This research additionally has four secondary contributions to the scientific community: a comparison of standard reinforcement learning to hierarchical reinforcement learning, an evaluation of two different ways to automatically create sub-goals, a comparison of manual and automatically created sub-goals, and a performance evaluation of the frequency of termination.
In this section, we discuss the results of the experiments and show why each experiment is important. Experiment one highlights how HRL can outperform traditional reinforcement learning in sparse reward environments. Experiments two and four together demonstrate the importance of sub-goals in HRL, as well as investigating different methods and quantities of sub-goals. Experiment three is essential as it shows the importance of having the algorithm choose the sub-goals.
@@ 6.1 Experiment 1: Comparison Between OC and PPO
OC was shown to perform better when compared to PPO as the complexity increases. The results show that with one-room multiple small obstacles maze, the PPO has a similar convergence performance to the OC, and as the maze complexity increases, the convergence performance of PPO compared to OC decreases steadily. This is shown in the Table 3 as the P-value for the one-room multiple small obstacles maze is 0.45, suggesting that the convergence performance is about the same. When the P-value for the one-room single large obstacle maze is less than 0.05 at 0.048, it is statistically significant, showing that the OC performs better than the PPO in terms of convergence time. With the four-rooms maze the difference is even larger where the OC converges about twice as fast as the PPO and has a P-value of 0.036. Although the convergence time is different between the PPO and OC, the path length is about the same. This was expected as they were both given time to converge, and they didn’t have difficulties finding a near-optimal maze. Based on these results, a T-test was not performed on the path lengths.
@@
Another conclusion we can draw from this experiment is OC’s ability to outperform PPO as the rewards become increasingly sparse. In the one-room multiple small obstacle maze, the reward is relatively easy to achieve as the obstacles are small and can be avoided relatively easily, but as the rewards get harder to reach in the one-room single large obstacle and the four-rooms maze PPO’s performance significantly decreases. This is because PPO can struggle as the rewards become increasingly sparse, but the OC is able to use the hierarchy it created to handle this sparsity.
@@ 6.3 Experiment 3: Comparison Between Automatic and Manual Sub-Goals
The idea behind sub-goals is to break down tasks into smaller tasks, such that the smaller tasks are easier for the policies to solve as they are not as complicated. But it is important that the task is split evenly. If not, one of the tasks will be more difficult and would remove the purpose of creating the smaller tasks to make easier tasks. Since it is important to break down the tasks correctly, it may seem best to set the sub-goals by hand to ensure the task is broken down correctly. To manually set the best possible sub-goals for this experiment, each doorway and the goal were chosen as the sub-goals. These were chosen because they evenly break down the task. This is similar to the originally designed four-rooms maze task, where each room was assigned a different option, with the doorway being the sub-goals SUTTON (40). As seen in Table 7 when the algorithm uses the manually set sub-goals it takes about twice as long to converge and it’s path length is about ten steps longer. These results support the hypothesis stating the algorithm will perform better if the sub-goals are set by the algorithm. These results support the hypothesis suggesting the algorithm is able to break down the tasks adequately, thus hand-set sub-goals are not required and even hinder the process as they impede exploration as most of the exploration in the Option Critic comes from trying different options and correspondingly pick a different action. The consideration is that if the sub-goals are set by hand, the algorithm wouldn’t be considering different options in the same state, explaining why automatic sub-goals produce better results. In summary the automatic sub-goals perform better because they allow the algorithm to explore more not because the sub-goals are necessarily better. In this research there was no metric to compare the sub-goals to ideal sub-goals but for future research such metric could be used to better understand the differences between sub-goals created by algorithm and sub-goals set by hand.
@@ 6.4 Experiment 4: Performance Evaluation on Frequency of Terminations
Theoretically, sub-goals are advantageous because they enable an algorithm to break down tasks into several smaller, simpler sub-tasks, which are easier for the algorithm to solve. These results support this hypothesis because the Option Critic is able to outperform the PPO as shown in Table 3. But Option Critics’ increased performance could come from other factors, in addition to sub-goals. In Table 8, it shows how as the termination regulation value increases from 0.00 to 0.10, the convergence time also increases. In Figure 12, it shows how as the termination regulation value increases, the sub-goals are longer and there are fewer sub-goals, up to the point where there are almost no sub-goals. In general, as the number of sub-goals decreases, the performance decreases, showing how it is important to have sub-goals. This might suggest that it would be best to terminate in every state, but as seen in experiment 3, this can also lead to poor performance.
6.5 Impact of Sub-goals and the Hierarchy
Current research into HRL has demonstrated its ability to outperform traditional reinforcement learning; however, there is a lack of research into why HRL performs better. Theoretically, it performs better due to its design to create a hierarchy for the task, but there is limited research on whether this is true in practice. Our research provides evidence of how sub-goals and building a hierarchy cause improvements in HRL.
Here, we demonstrate the importance of sub-goals by evaluating what happens when we remove them. There are two ways to remove the sub-goals: to remove the terminations and have one option or to terminate every step. The first option is essentially the same as traditional reinforcement learning, as a single option without terminations is a standard actor-critic policy. As discussed from the results of experiment one this results in worse performance. Additionally, in experiment four, we can see in Table 8 that as the termination regulation value increases, the number of sub-goals decreases, and as a result, the performance decreases. This shows why it is important to have at least some sub-goals but additional analysis shows how as the termination regulation value decreases the number of sub-goals increases and the performance increases as well. This suggests that the performance improvements are not coming from the sub-goals, but rather from terminating and switching options or policies. To test this we created experiment two, in experiment two we terminate every step to see if terminating the options is causing the performance increase. From experiment two’s results we can learn two things first the termination is not were the performance increase comes from, as terminating every step performs worse. The second is that even if we remove the termination functions and terminate every step instead; the critic will still create sub-goals. This explains why in experiment four the frequent terminations performed better because the best performance is when the termination function and the critic work together to create the sub-goals. Which is not possible if the options do not terminate frequently, as the critic is only used when the options are terminated. These experiments show the importance of sub-goals because if we remove them by not terminating or terminating every step then the performance decreases.
This illustrates the importance of sub-goals in most situations, but this may not be true in all cases. As there may be simple problems that do not need to be broken up and thus do not require sub-goals. Furthermore, in simple problems, sub-goals may hinder performance by adding unnecessary complexity to the task. By showing the importance of sub-goals, we also show the importance of the hierarchy since creating sub-goals creates a hierarchy as the sub-goals define when each sub-task ends and the next begins.
@@ 6.6 Future Work
Future research will investigate further ways to set the sub-goals automatically. Currently, the policy over options relies mostly on the critic to create sub-goals, but this could be replaced by using a different algorithm for the policy over options, such as an additional network or a different machine learning technique. This could improve the exploration of the sub-goals and the overall algorithm performance. Future research will also investigate the performance of the Option Critic in physical robot learning tasks in spatial navigation tasks in the context of sequence learning cazin (7) medeiros (23) Scleidorovich2022-zd (36).